OpenStudy (anonymous):

How would I factor this: 13x^2 - 13 = 0

OpenStudy (anonymous):

what is a common number found in the two terms 13x^2 and -13?

OpenStudy (anonymous):

13?

OpenStudy (solomonzelman):

yes!

OpenStudy (solomonzelman):

factor out of 13.

OpenStudy (anonymous):

13 and 1

OpenStudy (solomonzelman):

Well it is \[13x^2-13=0\]\[OR\]\[13 \times x^2 - 13 \times 1 = 0\]\[13(x^2-1)=0\]divide both sides by 13\[\frac{13(x^2-1)}{13}=\frac{0}{13}\]\[x^2-1=0\]\[Do ~~ it\]

OpenStudy (anonymous):

x=1

OpenStudy (solomonzelman):

or -1

OpenStudy (solomonzelman):

Yep! x=1 and x=-1